# Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?

### 20 Feb 2024 v2

@@ 4.2 Benchmarking QECC Variations
It is important to understand that the radar chart presented earlier highlights only the general comparative differences among various codes. We will now focus on surface codes, which are highly applicable in real-world scenarios. For illustration, let’s consider surface codes with distances of , , , and . Despite sharing common features like an identical error threshold, the same transversal gates, and equal access to decoding algorithms, these codes exhibit distinctions in several parameters. The structure of a surface code is a lattice, meaning the qubit overhead increases as the distance grows. All surface codes offer comparable error protection, which is enhanced with greater distances. Figure 11 (left) depicts the achievable logical error rates for varying physical error rates across different surface code distances. At a given physical error rate, a larger distance code yields a lower logical error rate. Moreover, all surface codes are scalable, but as the physical error rate rises, the necessity for greater distances becomes apparent. Although scalability is maintained with larger distances, it presents greater challenges. Figure 11 (right) illustrates the required distances to attain specific logical error rates for certain physical error rates. In practical terms, only the and distance surface codes have been implemented to date [9, 64, 43], with higher distances yet to be achieved. While the complexity of surface codes is generally considered high, it increases with distance. The graphs presented were generated through experiments conducted using STIM [32] simulations. Figure 12 presents a radar chart comparing four different surface codes against benchmarking parameters.
It is crucial to recognize that benchmarking is not only vital for evaluating different types of codes but also for assessing variations within existing codes. The process of selecting a code is dynamic and situational, with no fixed rules. Hence, paying attention to minor differences in codes is as important as considering major codes. In terms of code selection, although a universal solution doesn’t exist, we can establish a general hierarchy for benchmarking parameters based on typical priorities: The combined aspects of error protection and threshold typically receive the highest priority. This is followed by qubit overhead and complexity. Next in line are scalability and practical realization. Lastly, transversal gates and the number of decoding algorithms are considered. This ordering works because it aligns with the fundamental requirements of quantum computing, where error management is paramount, followed by resource efficiency (qubit usage and complexity) and practical feasibility (scalability and realization). Transversal gates and decoding algorithms, while important, are often tailored to specific use cases and thus are considered after these broader concerns. Figure 13 illustrates this prioritization of benchmarking parameters in code selection.
5 Future Work and Challanges
The field of quantum error correction is dynamic, continuously yielding diverse types of codes. This paper’s contribution — a set of parameters for benchmarking all QECCs — highlights an essential future direction: the evaluation of emerging codes. Understanding a code’s practicality and its position relative to existing solutions is crucial. Quantum error correction hinges on achieving a delicate balance; a comprehensive understanding of all parameters is vital for the appropriate application of each code.
Furthermore, the ongoing development in quantum error correction extends to existing codes and their variations. Over time, the realization of current QECCs across a broader spectrum of qubit technologies could alter their standings in the benchmark. While the parameters for evaluation remain constant, their relevance and implications may shift, underscoring the importance of continual updates and assessments in this rapidly evolving field.
6 Conclusion
This paper showcases the first universal benchmarking methodology for Quantum Error Correction Codes (QECCs), addressing a significant gap in quantum computing research. By meticulously evaluating eight QECCs against eight carefully selected parameters, this study not only establishes the importance of these parameters but also provides a nuanced understanding of how each QECC fares in various scenarios. The utilization of a radar chart for comparison vividly illustrates that every QECC involves distinct trade-offs, emphasizing that the selection of a QECC is a complex art of balancing diverse factors. This paper underlines the ongoing necessity of benchmarking as new and variations of QECCs are discovered, reflecting the dynamic nature of quantum computing where QECC selection is highly situational and no single solution prevails. This work is a step forward in ensuring that QECC choices are informed, balanced, and adaptable to evolving quantum computing needs.
Acknowledgements
The work is supported in parts by the National Science Foundation (NSF) (CNS-1722557, CCF-1718474, OIA-2040667, DGE-1723687, and DGE-1821766).

### 05 Mar 2024 v3

@@ IV-B Benchmarking QECC Variations
It is important to understand that the radar chart presented earlier highlights only the general comparative differences among various codes. We will now focus on surface codes, which are highly applicable in real-world scenarios. For illustration, let’s consider surface codes with distances of , , , and . Despite sharing common features like an identical error threshold, the same transversal gates, and equal access to decoding algorithms, these codes exhibit distinctions in several parameters. The structure of a surface code is a lattice, meaning the qubit overhead increases as the distance grows. All surface codes offer comparable error protection, which is enhanced with greater distances. Figure 11 (left) depicts the achievable logical error rates for varying physical error rates across different surface code distances [40]. At a given physical error rate, a larger distance code yields a lower logical error rate. Moreover, all surface codes are scalable, but as the physical error rate rises, the necessity for greater distances becomes apparent. Although scalability is maintained with larger distances, it presents greater challenges. Figure 11 (right) illustrates the required distances to attain specific logical error rates for certain physical error rates [40]. In practical terms, only the and distance surface codes have been implemented to date [80, 82, 17], with higher distances yet to be achieved. While the complexity of surface codes is generally considered high, it increases with distance. Figure 12 presents a radar chart comparing four different surface codes against benchmarking parameters. It is crucial to recognize that benchmarking is not only vital for evaluating different types of codes but also for assessing variations within existing codes.
V How to Strategically Choose a QECC?
V-A An Universal Streamlined Approach
The process of selecting a code is dynamic and situational, with no fixed rules. Hence, paying attention to minor differences in codes is as important as considering major codes. In terms of code selection, although a universal solution doesn’t exist, we can establish a general systematic strategy based on typical priorities. The selection strategy for QECCs, outlined in Fig. 13, involves a multi-stage filtration process, starting with an initial evaluation based on error types, qubit type, and feasibility. This is followed by a detailed examination of error thresholds and qubit overhead. The strategy emphasizes optimizing qubit overhead and scalability, revisiting previous stages if overhead is excessive or scalability is lacking. Complexity and transversal gates are then assessed to minimize both, while maximizing decoding algorithm options. The final QECC choice balances complexity against decoding efficiency, tailored to the scenario’s demands. A last review ensures the chosen QECC meets all criteria, with adjustments made to internal parameters as needed. If inadequacies are found, the process recommences.
V-B Introducing Our Tool for Strategic Recommendation
Our tool primarily aims to match the specific characteristics of a quantum computing scenario with suitable QECCs, recommending a list ranked from most to least favorable. This recommendation also includes the maximum feasible distance for each QECC, aligning with the understanding that longer distances typically correlate with lower logical error rates. Thus, our emphasis on maximizing distance is directly tied to achieving the lowest possible logical error rates. The framework evaluates scenarios based on five key factors: the qubit type being used, the available number of qubits for error correction within the system, the count of qubits in the circuit requiring correction, the presence of multi-qubit gates in the circuit, and the types and rates of physical errors encountered (including Pauli errors like bit-flip, phase-flip, or all-Pauli, as well as depolarizing, gate, and readout errors). Gate errors are the most weighted due to their extensive impact on the quantum system, followed by depolarizing and readout errors. Since all quantum errors can be categorized as Pauli errors, scenarios considering ‘all-Pauli’ errors are deemed most realistic. This flexibility aids in the versatility and characterization of QECCs. Each QECC includes a set of qubit types on which it has been implemented as one of its parameters. The ‘simulation’ type is consistently applied across all QECCs to accommodate those not yet realized in specific scenarios, ensuring that simulated implementations are considered for every QECC. The framework aims to optimize for reduced complexity in overall operations and lattice surgery, while enhancing the code distance and the availability of decoding algorithms. It generates a ranked list of QECCs from most to least suitable based on these considerations, along with the maximum achievable code distance for each QECC, determined by the user’s available qubits for error correction and the qubits in the original circuit. The fundamental operation of the framework is illustrated in Figure 14, providing a foundational overview.
The framework is designed to seamlessly integrate additional QECCs and to effortlessly update existing parameters. Whether a QECC is realized on a new qubit type previously unaccounted for, or if there’s an improvement in its error threshold, these changes can be rapidly incorporated. This adaptability ensures the framework remains current with advancements in the field. Moreover, the framework has an inbuilt ‘debug_mode’ which, when activated, allows users to observe the filtration process of QECCs step by step until only the final recommendations are left. Notably, the results are obtained almost instantaneously, highlighting the framework’s efficiency and speed. Table II illustrates the user inputs for three scenarios and the corresponding recommendations generated by the framework.
| Parameters | Scenario 1 | Scenario 2 | Scenario 3 | ||||||
| User Input | qType | supercond. | rydberg | simulation | |||||
| maxQAvail | 100 | 600 | 1500 | ||||||
| qOrig | 1 | 2 | 5 | ||||||
| multiQGate? | no | yes | no | ||||||
| errType | bit-flip | all-pauli | phase-flip | ||||||
| depErr | 1E -04 | 1E -04 | 1E -03 | ||||||
| gateErr | 1E -03 | 1E -03 | 1E -02 | ||||||
| readErr | 1E -02 | 1E -02 | 1E -01 | ||||||
| Recommendations |
HeavyH, 11 | ||||||
VI Conclusion
This paper presents the inaugural comprehensive benchmarking framework for Quantum Error Correcting Codes (QECCs), filling a critical void in the field of quantum computing research. Through a detailed comparison of eight QECCs across eight meticulously chosen parameters, our analysis not only highlights the pivotal role of these parameters but also offers an in-depth perspective on the performance of each QECC under different conditions. The innovative use of radar charts for illustrating comparisons clearly demonstrates the unique compromises associated with each QECC, underscoring the intricate process of QECC selection as an exercise in balancing a variety of factors. This study emphasizes the continuous need for benchmarking amidst the discovery of new QECCs and their variants, showcasing the fluid nature of quantum computing where the choice of QECC is highly contextual and no universal solution exists.
Furthermore, we propose a universal, structured methodology for selecting the most appropriate QECC for specific scenarios, aiding users in navigating the complex landscape of quantum computing to find a QECC that aligns with their unique needs. A significant achievement of this research is the creation of a flexible and dynamic pioneering tool that evaluates given scenario characteristics to recommend a prioritized list of QECCs, alongside the maximum achievable distance for each suggested code. This tool is built to accommodate future expansions, including the integration of new QECCs and the easy adjustment of their parameters, ensuring its ongoing utility and adaptability in the face of quantum computing’s rapid advancement. This work is a step forward in ensuring that QECC choices are informed, balanced, and adaptable to evolving quantum computing needs. 222GitHub Repository Link: Magic Mirror